OpenStudy (anonymous):

Which is the graph of y = 2.5x?

OpenStudy (perl):



OpenStudy (perl):

this goes through (0,0) and has slope 2.5

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

That is right I think.. :)

OpenStudy (anonymous):

What is the equation of the graphed line? A. y= -5/2x B. y= -2/5x C. y= 2/5x D. y= 5/2x

OpenStudy (anonymous):

Why are you asking your question now in reverse order? :P

OpenStudy (anonymous):

Tell me what is 5/2 = ?

OpenStudy (anonymous):

2 and a half

OpenStudy (anonymous):

Then what should be the option?

OpenStudy (anonymous):

c???

OpenStudy (anonymous):

Oh, 2/5 = ??

OpenStudy (anonymous):

is it c

OpenStudy (anonymous):

The graph you showed is of \(y = 2.5x\). now convert \(2.5\) to fraction.. :)

OpenStudy (anonymous):

5/2

OpenStudy (anonymous):

Then what is the correct option then?

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

As simple as that.. :)

OpenStudy (anonymous):

\(\text{This is it...}\)